OpenStudy (anonymous):

a local pet groomer must clip the nails of 40 birds and cats. if there are 110 feet, how many birds and cats are there?

OpenStudy (anonymous):

0 they all died i shot them

OpenStudy (cwrw238):

lol

OpenStudy (cwrw238):

x + y = 40 2x + 4y = 110 2x + 2y = 80 2y = 30 y = 15 there are 15 cats and 25 birds

OpenStudy (anonymous):

can i have a medal?

OpenStudy (anonymous):

There. you got a medal.